OpenStudy (anonymous):

how would you find the x an y intercepts of y=xsqrrt16-x^2 ?

OpenStudy (welshfella):

the y intercepts is when x = 0 so plug in x = 0 into the function and work it out

OpenStudy (welshfella):

to find the x intercepts sole xsqrrt16-x^2 = 0

OpenStudy (anonymous):

you will put y as equal to zero and then \[\sqrt{16}-x^{2}\] x will be 4 and -4

OpenStudy (welshfella):

i'm assuming there are parentheses around the 16 - x^2

OpenStudy (welshfella):

the x intercepts will be 4 and -4 but also 0.

OpenStudy (welshfella):

xsqrrt(16-x^2) = 0 x = 0 , x^2 - 16 = 0 x^2 = 16 so x = 4 ,-4

OpenStudy (welshfella):

y intercept = 0* sqrt16 = ?

OpenStudy (anonymous):

yint=0 ?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

thank you so much!

OpenStudy (welshfella):

yw